An express bus travels a certain 320km route in 1hr less time than it takes a local buss to travel a 300km route. Find the speed of each bus if the speed of the local is 20km/h less than that of the express.
Answer by sachi(548)

the speed of the local is 20km/h less than that of the express
so let the speed of the express bus=x & that of local=x-20
An express bus travels a certain 320km route in 1hr less time than it takes a local buss to travel a 300km route.
so time taken by both bussed are
the express took 320/x & the local took 300/(x-20)
as per question
300/(x-20)-320/x=1
or 300x-320(x-20)=x(x-20) multiplying x)x-20) on both sides
or 300x-320x+6400=x^2-20x
or -20x+6400=x^2-20x
or x^2=6400
or x=80 kmph= speed of express
& the speed of local=80-20=60 kmph